OpenStudy (anonymous):

Tickets to a show cost different for seniors, children, and adults. Adult tickets are twice as much as senior tickets. Two senior tickets plus one children’s ticket cost $17. Two children tickets plus one adult ticket cost $24. How much would it cost to buy one adult ticket and one senior ticket

OpenStudy (johnweldon1993):

Lets set up everything that we know Adult tickets cost twice as much as senior a = 2s two senior tickets plus 1 children's cost 17 2s + c = 17 Two children tickets plus one adult ticket cost $24 2c + a = 24 Now...How much would it cost to buy one adult ticket and one senior ticket? a = 2s 2s + c = 17 2c + a = 24 These are our 3 equations...how would we solve for 'a' and for 's'...?

OpenStudy (johnweldon1993):

Here I'll start you off...I'll solve for childrens tickets... Remember a = 2s...we can put this into our 3rd equation there 2c + 2s = 24 our second equation says that 2s + c = 17 Now lets look at that....2 senior tickets + 1 child = 17 but 2 seniors and 2 childrens = 24 This must mean that the price of a childs ticket.....would be the difference between 24 and 17....so a childs ticket would cost 7 dollars..... Now use that to solve for 's'